# On oscillator death in the Winfree model I: the one-dimensional case

### v1

@@ Abstract.
We show that for the standard sinusoidal Winfree model, a coupling strength exceeding twice the maximal magnitude of the intrinsic frequencies guarantees the convergence of the system for Lebesgue almost every initial data. This is proven by first showing, via an order parameter bootstrapping argument, that the pathwise critical coupling strength is upper bounded by a function of the order parameter, and then showing by a volumetric argument that for Lebesgue almost every data the order parameter cannot stay below and be bounded away from 1 for all time; this is a Winfree model counterpart of the analysis of Ha and the author (2020) performed for the Kuramoto model. Using concentration of measure and the aforementioned volumetric argument, we show that, except possibly on a set of very small measure, oscillator death is observed in finite time; this rigorously demonstrates the existence of the oscillator death regime numerically observed by Ariaratnam and Strogatz (2001). These results are robust under many other choices of interaction functions often considered for the Winfree model. We demonstrate that the asymptotic dynamics described in this paper are sharp by analyzing the equilibria of the Winfree model, and we bound the total number of equilibria using a polynomial description. The sequel to this paper will deal with the Winfree model on higher-dimensional manifolds, such as the Euclidean spheres and the unitary groups.
@@ 1. Introduction
The main result of this paper stated in the special case (2) is the following theorem.
@@
Numerical simulations suggest that the limiting equilibrium for almost every initial data is that of [HPR15], which is confined in a small neighborhood of 0 modulo $2\pi$, which becomes arbitrarily small as one takes $\kappa\to\infty$. We thus pose the following question.
Question 4.
@@
Remark 5.
On the other hand, [HPR15, Theorem 2.2] (stated below in Theorem 18) tells us that a variant of Question 4 where we interchange the order of quantifiers is affirmative:
@@
For each initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, denote by $\{\theta_{i}(t)\}_{i=1}^{N}$ the solution to (3). Then, denoting by $m$ the normalized Lebesgue measure on $\pi,\pi]^{N}$ such that $m([-\pi,\pi]^{N})=1$, one has
@@
$\qquad\forall i=1,\cdots,N~\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$
@@
$\geq 1-\left(\sqrt{\frac{\pi e}{2}}\frac{\|\Omega\|_{\infty}^{1/3}}{\kappa^{1/3}}\right)^{N}.$
@@
Fix parameters [^1] The reason we don’t consider the case $N=1$ is because oscillator death is trivial; see Proposition 46 with $N=1$. $N\geq 2$, $\{\omega_{i}\}_{i=1}^{N}\in\mathbb{R}^{N}$, $\kappa>0$, and $\varepsilon\in(0,1]$ such that
@@
$\qquad\forall i=1,\cdots,N~\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$
@@
Theorems 6 and 7 rigorously confirm the experimental observations of [AS01] in a more precise manner. Namely, Theorem 6 tells us that if $\kappa>(2+\varepsilon)\max_{i}|\omega_{i}|$, then if we independently sample $\theta_{i}^{0}$ uniformly from $\pi,\pi]$ and simulate (3), we will then observe that the oscillators stay within $2\pi$ of its initial value for all time with probability at least $1-\exp\left(-\frac{\varepsilon^{2}N}{25}\right)$, which tends rapidly to $1$ as $N\to\infty$. For example, if we have $N=800$ and $\kappa>3\max_{i}|\omega_{i}|$, then this probability is at least $1-\exp\left(-800/25\right)\approx 1-1.266\times 10^{-14}$. Theorem 7 goes one step further and tells us that oscillator death behavior will be observed within some large time $T$ with probability at least $1-\left(\frac{20}{\pi e\varepsilon}+\frac{4\kappa(5-\varepsilon)}{5\pi e}(T-T_{0})\right)^{-\frac{N}{2}}$ if $T$ is large enough. For example, if $N=800$, $\max_{i}|\omega_{i}|=2$, $\kappa=6$ (so that $\varepsilon=1$), and $T=500$, then $T_{0}=\frac{25}{4\kappa N}\log\left(\left(1+\frac{2}{N}\right)\left(\frac{2\sqrt{5}}{\sqrt{\pi}e^{1/2+1/25}}\right)^{N}-\frac{2}{N}\right)=0.40156699\cdots$ and the possibility of observing oscillator death within time $T=500$ is at least
@@
Many of the results of this paper can be extended to the manifold setting, particularly the Euclidean sphere $\mathbb{S}^{d-1}$ and the unitary group $\mathrm{U}(d)$. This will be dealt with in a sequel to this paper.
@@
In Section 5, we calculate the divergence of the vector field associated to system (1) and use an argument on the rate of change of volume to prove Theorem 1. By applying some large deviations theory and capitalizing on the positivity of the divergence, we will prove Theorems 6 and 7, plus many other new theorems mentioned in Section 2.
@@ 2.1. The pathwise critical coupling strength is bounded by a function of the order parameter
It is conceptually convenient to identify each $\theta_{j}$ with the corresponding point $e^{\mathrm{i}\theta_{j}}$ on the unit circle and consider (3) to be a dynamical system on $\mathbb{S}^{1}$. Then, the order parameter $R$ measures how far away and how close the points $e^{\mathrm{i}\theta_{j}}$ are from 1-1 and to $1$, respectively. Under this interpretation, Corollary 16 roughly says that if $\kappa\gtrsim{R_{0}}^{-3/2}\max_{i}|\omega_{i}|$, then (a) there is a uniform-in-time lower bound on the macroscopic quantity $R(t)$, (b) each particle $e^{\mathrm{i}\theta_{j}}$ will converge while not making a full revolution around the unit circle, and (c) any particle $e^{\mathrm{i}\theta_{j}}$ which is far away from 1-1 will be pulled towards and trapped near $1$. Thus, it is easier to constrain the behavior of (3) when $R_{0}$} is large, and when $R_{0}$} is small, we cannot say much about the long-term dynamics of (3) using Corollary 16 unless we pay the high price of $\kappa\gtrsim{R_{0}}^{-3/2}\max_{i}|\omega_{i}|$. Furthermore, if we are allowed to take $\kappa$ to be arbitrarily large, then (f) we can guarantee $R(t)$ to not go below its initial value by any prescribed amount, and (g) for any neighborhood of 1-1 in $\mathbb{S}^{1}$ we may bring any oscillator not in that neighborhood close to $1$ with arbitrary accuracy. All this happens while (e,i) the oscillators $e^{\mathrm{i}\theta_{j}}$ close to 1 are linearly ordered according to their natural frequencies $\omega_{j}$.
@@
Although the statement of Theorem 6 treats $2\max_{i}|\omega_{i}|$ as if it were a critical threshold for $\kappa$, due to the above argument of inserting $R_{0}=1$ into Corollary 16, this is likely an artifact of our proof, as the phase diagram of [AS01] indicates that complete oscillator death happens even for smaller values of $\kappa$. The value $2\max_{i}|\omega_{i}|$ for $\kappa$ comes from the case $R_{0}=1$ of (8), but note that in the endpoint case $R_{0}=2$ of (8), the value $\frac{4}{3\sqrt{3}}\|\Omega\|_{\infty}$ for $\kappa$ is the maximum of $\kappa_{c}(\Omega)$ ranging over all $\Omega$ with a fixed $\|\Omega\|_{\infty}$. In other words, for a fixed $\|\Omega\|_{\infty}=\max_{i}|\omega_{i}|$, $\kappa_{\mathrm{crit}}\coloneqq\frac{4}{3\sqrt{3}}\|\Omega\|_{\infty}$ is a critical threshold for the existence of equilibria in the sense that if $\kappa>\kappa_{\mathrm{crit}}$ then (3), by [HPR15] (see Theorem 18 below), has an asymptotically stable equilibrium (i.e., an equilibrium with an open neighborhood such that for any initial data in that open set the solution to (3) converges to that equilibrium), while if $|\kappa|<\kappa_{\mathrm{crit}}$ and $\omega_{1}=\cdots=\omega_{N}=\|\Omega\|_{\infty}$ then, by [HKPR17, OKT17] not only does (3) fail to have any equilibria, but there is a positive measure of solutions to (3) satisfying $\theta_{i}(t)\to\infty$ as $t\to\infty$, $i=1,\cdots,N$.
Here is the theorem mentioned above.
@@
    (Exponential convergence to equilibrium) we have, for $q\geq 1$, that
    $\|\Theta(t)-\Theta^{\infty}\|_{\ell_{q}^{N}}\leq\|\Theta(T)-\Theta^{\infty}\|_{\ell_{q}^{N}}\exp\left[-\kappa(2\cos\alpha^{\infty}-1)(\cos\alpha^{\infty}+1)(t-T)\right],\quad t\geq T.$
@@
2.2. The Łojasiewicz gradient theorem
@@ 2.3. General interaction functions
Under stronger conditions on $I$ and $S$, we can take advantage of the time flow and can deduce the following uniform oscillator death result which is an analogue of Theorem 1.
@@ 5.1. Some large deviations theory
$\qquad\forall i=1,\cdots,N~\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$
@@ 5.2.1. The prototypical Winfree model (3)
$\qquad\forall i=1,\cdots,N~\exists\lim_{t\to\infty}\theta_{i}(t),~\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi,~\lim_{t\to\infty}\dot{\theta}_{i}(t)=0$

### v2

@@ Abstract.
We show that for the standard sinusoidal Winfree model, a coupling strength exceeding twice the maximal magnitude of the intrinsic frequencies guarantees the convergence of the system for Lebesgue almost every initial data. This is proven by first showing, via an order parameter bootstrapping argument, that the pathwise critical coupling strength is upper bounded by a function of the order parameter, and then showing by a volumetric argument that for Lebesgue almost every data the order parameter cannot stay below and be bounded away from 1 for all time; this is a Winfree model counterpart of the analysis of Ha and the author (2020) performed for the Kuramoto model. Using concentration of measure and the aforementioned volumetric argument, we show that, except possibly on a set of very small measure, oscillator death is observed in finite time; this rigorously demonstrates the existence of the oscillator death regime numerically observed by Ariaratnam and Strogatz (2001). These results are robust under many other choices of interaction functions often considered for the Winfree model. We demonstrate that the asymptotic dynamics described in this paper are sharp by analyzing the equilibria of the Winfree model, and we bound the total number of equilibria using a polynomial description.
@@ 1. Introduction
The main result of this paper stated in the special case (2) is the following theorem. For more general interaction functions, see subsection 2.3 for analogous results.
@@
Remark 4.
On the other hand, [HPR15, Theorem 2.2] (stated below in Theorem 18) gives an equilibrium confined in a small neighborhood of 0 modulo $2\pi$, which becomes arbitrarily small as one takes $\kappa\to\infty$, and whose basin of attraction covers almost all of the torus as one takes $\kappa\to\infty$. This tells us that a variant of Question 5 where we interchange the order of quantifiers is affirmative:
Numerical simulations suggest that the limiting equilibrium for almost every initial data is that of [HPR15]. We thus pose the following question.
Question 5.
@@
For each initial data $\{\theta_{i}^{0}\}_{i=1}^{N}$, denote by $\{\theta_{i}(t)\}_{i=1}^{N}$ the solution to (3). Then, denoting by $m$ the normalized Lebesgue measure on $\pi,\pi]^{N}$ such that $m([-\pi,\pi]^{N})=1$, one has [^1] By Theorem 1 we may harmlessly insert the condition $\exists\lim_{t\to\infty}\theta_{i}(t)$ and $\lim_{t\to\infty}\dot{\theta}_{i}(t)$ for all $i=1,\cdots,N$” into the events below.
@@
$\qquad\forall i=1,\cdots,N~\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,$
@@
$m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi~\mbox{for all }i=1,\cdots,N\right\}\geq 1-\left(\sqrt{\frac{\pi e}{2}}\frac{\|\Omega\|_{\infty}^{1/3}}{\kappa^{1/3}}\right)^{N}.$
@@
Fix parameters [^2] The reason we don’t consider the case $N=1$ is because oscillator death is trivial in this case; see Proposition 46 with $N=1$. $N\geq 2$, $\{\omega_{i}\}_{i=1}^{N}\in\mathbb{R}^{N}$, $\kappa>0$, and $\varepsilon\in(0,1]$ such that
@@
$\qquad\forall i=1,\cdots,N~\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi,$
@@
$m\left\{\{\theta_{i}^{0}\}_{i=1}^{N}\in[-\pi,\pi]^{N}:\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi~\mbox{for all }i=1,\cdots,N\right\}$
@@
Theorems 6 and 7 rigorously confirm the experimental observations of [AS01]. Namely, Theorem 6 tells us that if $\kappa>(2+\varepsilon)\max_{i}|\omega_{i}|$, then if we independently sample $\theta_{i}^{0}$ uniformly from $\pi,\pi]$ and simulate (3), we will then observe that the oscillators stay within $2\pi$ of its initial value for all time with probability at least $1-\exp\left(-\frac{\varepsilon^{2}N}{25}\right)$, which tends rapidly to $1$ as $N\to\infty$. For example, if we have $N=800$ and $\kappa>3\max_{i}|\omega_{i}|$, then this probability is at least $1-\exp\left(-800/25\right)\approx 1-1.266\times 10^{-14}$. Theorem 7 goes one step further and tells us that oscillator death behavior will be observed within some large time $T$ with probability at least $1-\left(\frac{20}{\pi e\varepsilon}+\frac{4\kappa(5-\varepsilon)}{5\pi e}(T-T_{0})\right)^{-\frac{N}{2}}$ if $T$ is large enough. For example, if $N=800$, $\max_{i}|\omega_{i}|=2$, $\kappa=6$ (so that $\varepsilon=1$), and $T=500$, then $T_{0}=\frac{25}{4\kappa N}\log\left(\left(1+\frac{2}{N}\right)\left(\frac{2\sqrt{5}}{\sqrt{\pi}e^{1/2+1/25}}\right)^{N}-\frac{2}{N}\right)=0.40156699\cdots$ and the possibility of observing oscillator death within time $T=500$ is at least
@@
The results stated above extend to a larger class of interaction functions $I$ and $S$ for (1); see subsection 2.3 for the statements. Also, many of the results of this paper can be extended to the manifold setting, particularly the Euclidean sphere $\mathbb{S}^{d-1}$ and the unitary group $\mathrm{U}(d)$. We will deal with this in a sequel to this paper.
@@
In Section 5, we calculate the divergence of the vector field associated to system (1) and use an argument on the rate of change of volume to prove Theorem 1. By applying some large deviations theory and capitalizing on the positivity of the divergence, we will prove Theorems 6 and 7, plus the many other new theorems mentioned in Section 2.
@@ 2.1. The pathwise critical coupling strength is bounded by a function of the order parameter
It is conceptually convenient to identify each $\theta_{j}$ with the corresponding point $e^{\mathrm{i}\theta_{j}}$ on the unit circle and consider (3) to be a dynamical system on the $N$-torus $\mathbb{T}^{N}=(\mathbb{S}^{1})^{N}$. Then, the order parameter $R$ measures how far away and how close the points $e^{\mathrm{i}\theta_{j}}$ are from 1-1 and to $1$, respectively. Under this interpretation, Corollary 16 roughly says that if $\kappa\gtrsim{R_{0}}^{-3/2}\max_{i}|\omega_{i}|$, then (a) there is a uniform-in-time lower bound on the macroscopic quantity $R(t)$, (b) each particle $e^{\mathrm{i}\theta_{j}}$ will converge while not making a full revolution around the unit circle, and (c) any particle $e^{\mathrm{i}\theta_{j}}$ which is far away from 1-1 will be pulled towards and trapped near $1$. Thus, it is easier to constrain the behavior of (3) when $R_{0}$} is large, and when $R_{0}$} is small, we cannot say much about the long-term dynamics of (3) using Corollary 16 unless we pay the high price of $\kappa\gtrsim{R_{0}}^{-3/2}\max_{i}|\omega_{i}|$. Furthermore, if we are allowed to take $\kappa$ to be arbitrarily large, then (f) we can guarantee $R(t)$ to not go below its initial value by any prescribed amount, and (g) for any neighborhood of 1-1 in $\mathbb{S}^{1}$ we may bring any oscillator not in that neighborhood close to $1$ with arbitrary accuracy. All this happens while (e,i) the oscillators $e^{\mathrm{i}\theta_{j}}$ close to 1 become linearly ordered according to their natural frequencies $\omega_{j}$.
@@
Although the statement of Theorem 6 treats $2\max_{i}|\omega_{i}|$ as if it were a critical threshold for $\kappa$, due to the above argument of inserting $R_{0}=1$ into Corollary 16, this is likely an artifact of our proof, as the phase diagram of [AS01] indicates that complete oscillator death happens even for smaller values of $\kappa$. The value $2\max_{i}|\omega_{i}|$ for $\kappa$ comes from the case $R_{0}=1$ of (8), but note that in the endpoint case $R_{0}=2$ of (8), the value $\frac{4}{3\sqrt{3}}\|\Omega\|_{\infty}$ for $\kappa$ is the maximum of $\kappa_{c}(\Omega)$ ranging over all $\Omega$ with a fixed $\|\Omega\|_{\infty}$. In other words, for a fixed $\|\Omega\|_{\infty}=\max_{i}|\omega_{i}|$, $\kappa_{\mathrm{crit}}\coloneqq\frac{4}{3\sqrt{3}}\|\Omega\|_{\infty}$ is a another critical threshold for the existence of equilibria in the sense that if $\kappa>\kappa_{\mathrm{crit}}$ then (3), by [HPR15] (see Theorem 18 below), has an asymptotically stable equilibrium (i.e., an equilibrium with an open neighborhood such that for any initial data in that open set the solution to (3) converges to that equilibrium), while if $|\kappa|<\kappa_{\mathrm{crit}}$ and $\omega_{1}=\cdots=\omega_{N}=\|\Omega\|_{\infty}$ then, by [HKPR17, OKT17] not only does (3) fail to have any equilibria, but there is a positive measure of solutions to (3) satisfying $\theta_{i}(t)\to\infty$ as $t\to\infty$, $i=1,\cdots,N$.
Here is the theorem of [HPR15] mentioned in the previous paragraph.
@@
    (Exponential convergence to equilibrium) we have that
    $\|\Theta(t)-\Theta^{\infty}\|_{\ell_{1}^{N}}\leq\|\Theta(T)-\Theta^{\infty}\|_{\ell_{1}^{N}}\exp\left[-\kappa(2\cos\alpha^{\infty}-1)(\cos\alpha^{\infty}+1)(t-T)\right],\quad t\geq T.$
@@
2.2. Application of the Łojasiewicz gradient theorem
@@ 2.3. General interaction functions
Under different conditions on $I$ and $S$, we can take advantage of the time flow and can deduce the following uniform oscillator death result which is an analogue of Theorem 1.
@@ 5.1. Some large deviations theory
$\qquad\forall i=1,\cdots,N~\sup_{t\geq 0}\theta_{i}(t)-\inf_{t\geq 0}\theta_{i}(t)<2\pi,$
@@ 5.2.1. The prototypical Winfree model (3)
$\qquad\forall i=1,\cdots,N~\sup_{t\geq T}\theta_{i}(t)-\inf_{t\geq T}\theta_{i}(t)<2\pi,$